Fine the point- slope equation of the line containing (-3, -4) and having slope m = -2.
Answer by t0hierry(194)

(y - (-4 ))= -2(x -(-3))
Simplifying
y+4 = -2(x + 3) or
y = -2x -10
Alternatively
y = -2(x+5)
check:
-4 = -2(-3 + 5) is this true?
4 =2(-3+5)
yes.